# An Interdisciplinary and Cross-Task Review on Missing Data Imputation

### v1

@@ 5.3 Hot-Deck Imputation
Hot-deck imputation (Andridge and Little, 2010) replaces the missing values of one or several variables for a non-respondent (known as the recipient) with observed values from a respondent (referred to as the donor) who shares similar characteristics with the non-respondent. Some variations of this method involve randomly selecting the donor from a group of potential donors, also known as the donor pool. These are termed as random hot deck methods. Conversely, some versions identify a single donor, typically the nearest neighbor (NN) based on a certain metric, and impute values from that particular case. These are referred to as deterministic hot-deck methods, as they don’t involve any randomness in the selection of the donor. There are also methods that impute summaries of values from a group of donors, like the mean or weighted average, instead of individual values. The process of hot-deck imputation typically involves the following steps:
@@
This method generally offers greater accuracy than basic imputation strategies, as it replaces missing data with plausible values derived from similar, observed cases rather than a single statistic like the mean, which can distort the underlying variable distributions. A key advantage is that the imputed values preserve the original data’s distributional characteristics because they are conditioned on the matching (categorization) variables. This approach is particularly beneficial when data is missing in specific patterns (Roth, 1994). The methodology, implementation, and applications of this technique, known as hot-deck imputation, are detailed in (Andridge and Little, 2010).
For hot-deck imputation, the nearest neighbors or the donor pool are found by using a distance or similarity metric. There are many choices for the distance or similarity metric. For numerical data, one may use the Euclidean distance, $\ell_{1}$ distance, cosine similarity, Mahalanobis distance, Chebyshev distance, etc (Fkih, 2022). For categorical data, one may use the overlap measure (or Hamming distance), Eskin (Eskin et al., 2002), Goodall’s measures (Goodall, 1966), Gambaryan measure, inverse occurrence frequency, etc. A review of similarity measures for categorical data can be found in (Chandola et al., 2007). Consequently, for mixed-type data, one needs to combine different distance or similarity measures.
The most representative example of hot-deck imputation is the kNN imputation (Chen and Shao, 2000; Beretta and Santaniello, 2016), which is a deterministic hot-deck method. Suppose ${x}_{ij}$, the variable $j$ at observation $i$, is a missing value. kNN imputation predicts $\hat{x}_{ij}$ by the following weighted sum
@@
where $\mathcal{N}_{k}$ denotes the index set of the $k$ nearest neighbors of observation $i$ without missing values of variable $j$. $w_{v}$ are the weights of the neighbors and satisfy $\sum_{v\in\mathcal{N}_{k}}w_{v}=1$. If $w_{v}$, $v\in\mathcal{N}_{k}$, are all $1/k$, $\hat{x}_{ij}$ is just the average of the $k$ nearest neighbors. In recommendation systems, $w_{v}$ are usually set to $w_{v}=s_{vi}/\sum_{l\in\mathcal{N}_{k}}s_{li}$ (Breese et al., 1998; Sarwar et al., 2001), where $s_{li}$ denotes the similarity between the observations $l$ and $i$. Note that $s_{li}$ is computed from only the overlapped observed values of the two samples.
Hot-deck imputation is widely used due to its relative simplicity and its non-parametric nature, as it does not require assumptions about the underlying data distribution. However, the method has several limitations. It can underestimate data variability by redistributing existing values without introducing new information. Furthermore, imputation quality is highly dependent on the chosen similarity metric; a poor metric that fails to capture true associations between cases can lead to inaccurate imputations. This issue is exacerbated with high missingness rates or noisy data.
@@ 5.4 Likelihood Methods
Maximum likelihood is a strategy for estimating the unknown parameters of a model from data. When data is complete, it’s straightforward to calculate quantities like the mean, covariance, and linear regression coefficients, which are maximum likelihood estimates derived from maximizing the likelihood of the observed data. This same principle applies even when there are missing values in the data, as the estimation can still be based on the likelihood of the observed values. However, when missing data is present, the likelihood of the observed data becomes more complex compared to the typical data analysis scenarios, and finding an estimate that maximizes this likelihood doesn’t have a straightforward solution. To tackle the challenge, Dempster et al. (1977) proposed the EM algorithm for incomplete data. The main idea and algorithm are explained in the following context.
Following the notations defined by Section 3.1 and letting $\theta$ and $\phi$ be the parameters of the data model and the parameters of the missing mechanism respectively, the full likelihood can be formulated as
@@
The two steps are repeated until some convergence conditions are met. Finally, the missing values are filled by using the estimated $\theta$. See the example of multivariate Gaussian in Section 11.6 of (Murphy, 2012).
Despite its effectiveness, the Expectation-Maximization (EM) algorithm has several major limitations. First, it relies on the Missing at Random (MAR) assumption and is not suitable for data that are Missing Not at Random (MNAR). Second, its performance is contingent on a pre-specified data distribution (e.g., Gaussian); violations of this assumption can lead to inaccurate imputations. Finally, the algorithm is computationally intensive, and its results can be sensitive to the initial parameter values.
@@ 5.5.1 Low-Rank Matrix Completion
Low-rank matrix completion (LRMC) (Srebro and Shraibman, 2005; Candès and Recht, 2009; Wen et al., 2012; Hardt, 2014) addresses the problem of recovering missing entries in a matrix under the assumption that the true matrix is low-rank. Formally, a matrix is low-rank if its rank (the dimension of the vector space spanned by its rows or columns) is significantly smaller than its dimensions. An equivalent characterization is that the matrix has a rapidly decaying singular value spectrum, with only a few non-zero or significant singular values. Consider a matrix $\mathbf{X}\in\mathbb{R}^{m\times n}$, where $m<n$, the singular value decomposition (SVD) is denoted as $\mathbf{X}=\mathbf{U}\bm{\Sigma}\mathbf{V}^{\top}$, where $\mathbf{U}\in\mathbb{R}^{m\times m}$ and $\mathbf{V}\in\mathbb{R}^{n\times n}$ are orthonormal matrices, i.e., $\mathbf{U}^{\top}\mathbf{U}=\mathbf{I}_{m}$ and $\mathbf{V}^{\top}\mathbf{V}=\mathbf{I}_{n}$, $\bm{\Sigma}=(\text{diag}(\sigma_{1},\sigma_{2},\ldots,\sigma_{m}),\bm{0})$ and $\sigma_{1}\geq\sigma_{2}\geq\cdots\geq\sigma_{m}$ are the singular values. The rank of $\mathbf{X}$ is defined as $\text{rank}(\mathbf{X}):=\sum_{i=1}^{m}\mathbbm{1}(\sigma_{i}\neq 0)$. It has been found that many data matrices in real applications are low-rank ($\text{rank}(\mathbf{X})\ll m$) or can be well approximated by a low-rank matrix, i.e., many singular values are close to zero.
In the past decades, many LRMC methods have been proposed. They can be categorized into two groups. The first group is low-rank regularization methods, and the other group is low-rank factorization methods. Given an incomplete data matrix $\mathbf{X}\in\mathbb{R}^{m\times n}$, where the set of locations of observed entries is denoted as $\Omega$, the low-rank regularization methods aim to solve the following problem
[…]
subject to $\mathcal{P}_{\Omega}(\hat{\mathbf{X}})=\mathcal{P}_{\Omega}(\mathbf{X}),$
@@
subject to $\mathcal{P}_{\Omega}(\hat{\mathbf{X}})=\mathcal{P}_{\Omega}(\mathbf{X})$
@@
$\underset{\mathbf{A},\mathbf{B}}{\operatorname{minimize}}$ $\|\mathcal{P}_{\Omega}(\mathbf{Y}-\mathbf{AB})\|_{F}^{2}+\lambda\cdot R(\mathbf{A},\mathbf{B})$ (10)
where $\lambda>0$ is a hyperparameter and $R$ denotes the regularization on the factors $\mathbf{A}$ and $\mathbf{B}$. Table 3 summarizes the choices for $R$, where most of them are closely related to the nuclear norm and Schatten-$p$ quasi-norms. These regularizers are more efficient to implement than those in Table 2 because they do not require SVD or only need to perform SVD on the smaller factors. Once $\mathbf{A}$ and $\mathbf{B}$ are found, the recovered matrix is $\hat{\mathbf{X}}=\mathbf{A}\mathbf{B}$.
@@
nuclear norm $\|\mathbf{X}\|_{\ast}=\sum_{i=1}^{m}\sigma_{i}(\mathbf{X})$ (Candès and Recht, 2009)
truncated nuclear norm $\|\mathbf{X}\|_{\text{TN}}=\sum_{i=r+1}^{m}\sigma_{i}(\mathbf{X})$ (Hu et al., 2012)
weighted nuclear norm $\|\mathbf{X}\|_{\text{WN}}=\sum_{i=1}^{m}w_{i}\sigma_{i}(\mathbf{X})$ (Gu et al., 2014)
Schatten-$p$ quasi-norm $\|\mathbf{X}\|_{\text{Sp}}=(\sum_{i=1}^{m}\sigma_{i}^{p}(\mathbf{X}))^{1/p}$ (Nie et al., 2012)
weighted Schatten-$p$ quasi-norm $\|\mathbf{X}\|_{\text{WSp}}=(\sum_{i=1}^{m}w_{i}\sigma_{i}^{p}(\mathbf{X}))^{1/p}$ (Xie et al., 2016b)
generalized non-convex penalty $R(\mathbf{X})=\sum_{i=1}^{m}g_{\lambda}(\sigma_{i}(\mathbf{X}))$ (Lu et al., 2014)
@@
factored nuclear norm $R(\mathbf{A},\mathbf{B})=\frac{1}{2}(\|\mathbf{A}\|_{F}^{2}+\|\mathbf{B}\|_{F}^{2})$ $\min R(\mathbf{A},\mathbf{B})=\|\mathbf{X}\|_{\ast}$ (Srebro et al., 2005)
max norm $R(\mathbf{A},\mathbf{B})=(\max_{i}\left\|\mathbf{A}_{i:}\right\|_{2})(\max_{j}\left\|\mathbf{B}_{:j}\right\|_{2})$ Srebro and Shraibman (2005)
factored Schatten-$1/2$ $R(\mathbf{A},\mathbf{B})=\frac{1}{2}(\|\mathbf{A}\|_{*}+\|\mathbf{B}\|_{*})$ $\min R^{2}(\mathbf{A},\mathbf{B})=\|\mathbf{X}\|_{\mathrm{S}_{1/2}}$ (Shang et al., 2016)
factored Schatten-$1/2$ $R(\mathbf{A},\mathbf{B})=\frac{1}{2}(\|\mathbf{A}\|_{2,1}+\left\|\mathbf{B}^{T}\right\|_{2,1})$ $\min R^{2}(\mathbf{A},\mathbf{B})=\|\mathbf{X}\|_{\mathrm{S}_{1/2}}$ (Fan et al., 2019)
factored Schatten-$2q/(2+q)$ $R(\mathbf{A},\mathbf{B})=\frac{1}{q}\|\mathbf{A}\|_{2,q}^{q}+\frac{\alpha}{2}\left\|\mathbf{B}^{T}\right\|_{F}^{2}$ $\min R(\mathbf{A},\mathbf{B})=c\|\mathbf{X}\|_{S_{2q/(2+q)}}^{2q/(2+q)}$ (Fan et al., 2019)
factored Schatten-$p$ $R(\mathbf{A},\mathbf{B})=\frac{1}{2^{p}}\sum_{i=1}^{d}\left(\left\|\mathbf{A}_{:i}\right\|_{2}^{2}+\left\|\mathbf{B}_{i:}\right\|_{2}^{2}\right)^{p}$ $\min R(\mathbf{A},\mathbf{B})=\|\mathbf{X}\|_{S_{p}}^{p}$ (Giampouras et al., 2020)
@@
Beyond general low-rank matrices, several LRMC methods are designed for specific types. Notable examples include distance matrix completion (Alfakih et al., 1999), covariance matrix completion (Hippert-Ferrer et al., 2022), kernel matrix completion (Tsuda et al., 2003), and one-bit matrix completion (Davenport et al., 2014). In the latter case, the observed data are binary (1-bit) measurements rather than precise numerical values. This approach is valuable in scenarios with highly quantized data or binary feedback, such as in recommender systems and social networks.
@@ 5.5.2 High-Rank Matrix Completion
where $f^{(1)},\ldots,f^{(s)}$ are different nonlinear functions. This forms a matrix $\mathbf{X}=[\mathbf{X}^{(1)},\mathbf{X}^{(2)},\ldots,\mathbf{X}^{(s)}]$ of size $m\times n$, where $\mathbf{X}^{(j)}=[\mathbf{x}_{1}^{(j)},\ldots,\mathbf{x}^{(j)}_{n_{j}}]\in\mathbb{R}^{m\times n_{j}}$ and $n=\sum_{j=1}^{s}n_{j}$. It can be seen that (13) degrades to (12) if $s=1$ and degrades to (11) if all $f^{(j)}$ are linear. Figure 8 provides synthetic examples for (11),(12), and (13), respectively. The completion of the matrix for the data given by (13) is the most difficult.
Figure 8: Examples (Fan et al., 2020a) of data forming high-rank matrices in 3D space (left: union of subspaces; middle: one nonlinear manifold; right: union of nonlinear manifolds).
In recent years, a few high-rank matrix completion (HRMC) methods have been proposed (Eriksson et al., 2011; Alameda-Pineda et al., 2016; Elhamifar, 2016; Fan and Cheng, 2018; Fan et al., 2020a; Le Morvan et al., 2020). They can be organized into three categories: subspace-based methods, kernel-based methods, and deep learning-based methods.
@@
These methods were proposed for the data given by (11) and usually learn the subspaces from the incomplete data (Eriksson et al., 2011; Fan et al., 2020b) or exploit the self-expressive model (Yang et al., 2015; Elhamifar, 2016; Fan and Chow, 2017c, b), where each data point can be well represented by other data points drawn from the same subspace. The self-expressive model (Elhamifar and Vidal, 2013) can be formulated as
@@
where $\mathbf{S}\in\mathbb{R}^{n\times n}$ is the coefficient matrix and is usually assumed to be sparse. Suppose $\mathbf{X}$ is incomplete, Fan and Chow (2017b) proposed the following matrix completion method:
@@
subject to $~\mathcal{P}_{\Omega}(\hat{\mathbf{X}})=\mathcal{P}_{\Omega}(\mathbf{X})$
@@
These methods (Alameda-Pineda et al., 2016; Ongie et al., 2017; Fan and Chow, 2018; Fan and Udell, 2019b; Fan et al., 2020a; Ongie et al., 2021) take advantage of kernel techniques that can convert a nonlinear problem in the original data space to a linear problem in the high-dimensional feature space. Ongie et al. (2017) and Fan and Chow (2018) proposed to solve
@@
subject to $~\mathcal{P}_{\Omega}(\hat{\mathbf{X}})=\mathcal{P}_{\Omega}(\mathbf{X})$
where $\mathbf{K}\in\mathbb{R}^{n\times n}$ is a kernel matrix with $K_{i,j}=k(\mathbf{x}_{i},\mathbf{x}_{j})$ and $k$ is a kernel function such as the Gaussian kernel $k(\mathbf{x}_{i},\mathbf{x}_{j})=\exp(-\gamma\|\mathbf{x}_{i}-\mathbf{x}_{j}\|^{2})$. The motivation is that the feature matrix $\phi(\mathbf{X})$ induced by $k$ is approximately low-rank when $\mathbf{X}$ has some nonlinear latent structure and $\|\phi(\mathbf{X})\|_{S_{p}}^{p}=\text{trace}((\phi^{\top}(\mathbf{X})\phi(\mathbf{X}))^{p/2})=\text{trace}(\mathbf{K}^{p/2})$.
Fan and Udell (2019b) showed that $\text{rank}(\phi(\mathbf{X}))=\min\{\binom{m+q}{q},n,s\binom{d+pq}{pq}\}$ if $f^{(j)}$ are $p$-order polynomials and $\phi$ is the feature map of a $q$-order polynomial kernel. When $n$ is sufficiently large and $p$ and $q$ are relatively small, the rank of $\phi(\mathbf{X})$ is low compared to its dimensions. Fan and Udell (2019b) introduced a kernel factorization method for HRMC and an online algorithm, which are more efficient than the methods proposed in (Ongie et al., 2017; Fan and Chow, 2018).
@@
These methods (Fan and Chow, 2017a; Fan and Cheng, 2018; Seyedi et al., 2023) are based on neural networks and will be detailed in Section 5.6.

### v2

@@ 5.3 Hot-Deck Imputation
Hot-deck imputation (andridge2010review) replaces the missing values of one or several variables for a non-respondent (known as the recipient) with observed values from a respondent (referred to as the donor) who shares similar characteristics with the non-respondent. Some variations of this method involve randomly selecting the donor from a group of potential donors, also known as the donor pool. These are termed as random hot deck methods. Conversely, some versions identify a single donor, typically the nearest neighbor (NN) based on a certain metric, and impute values from that particular case. These are referred to as deterministic hot-deck methods, as they don’t involve any randomness in the selection of the donor. There are also methods that impute summaries of values from a group of donors, like the mean or weighted average, instead of individual values. The process of hot-deck imputation typically involves the following steps:
@@
This method generally offers greater accuracy than basic imputation strategies, as it replaces missing data with plausible values derived from similar, observed cases rather than a single statistic like the mean, which can distort the underlying variable distributions. A key advantage is that the imputed values preserve the original data’s distributional characteristics because they are conditioned on the matching (categorization) variables. This approach is particularly beneficial when data is missing in specific patterns (roth1994missing). The methodology, implementation, and applications of this technique, known as hot-deck imputation, are detailed in (andridge2010review).
For hot-deck imputation, the nearest neighbors or the donor pool are found by using a distance or similarity metric. There are many choices for the distance or similarity metric. For numerical data, one may use the Euclidean distance, $\ell_{1}$ distance, cosine similarity, Mahalanobis distance, Chebyshev distance, etc (FKIH20227645). For categorical data, one may use the overlap measure (or Hamming distance), Eskin (eskin2002geometric), Goodall’s measures (goodall1966new), Gambaryan measure, inverse occurrence frequency, etc. A review of similarity measures for categorical data can be found in (chandola2007similarity). Consequently, for mixed-type data, one needs to combine different distance or similarity measures.
The most representative example of hot-deck imputation is kNN imputation (chen2000nearest; beretta2016nearest), which is a deterministic hot-deck method. Suppose ${x}_{ij}$, the variable $j$ at observation $i$, is a missing value. kNN imputation make prediction for ${x}_{ij}$ by the following weighted sum
@@
where $\mathcal{N}_{k}$ denotes the index set of the $k$ nearest neighbors of observation $i$ without missing values of variable $j$. $w_{v}$ are the weights of the neighbors and satisfy $\sum_{v\in\mathcal{N}_{k}}w_{v}=1$. If $w_{v}$, $v\in\mathcal{N}_{k}$, are all $1/k$, $\hat{x}_{ij}$ is just the average of the $k$ nearest neighbors. In recommendation systems, $w_{v}$ are usually set to $w_{v}=s_{vi}/\sum_{l\in\mathcal{N}_{k}}s_{li}$ (breeseCF1998; sarwar2001item), where $s_{li}$ denotes the similarity between observations $l$ and $i$. Note that $s_{li}$ is computed from only the overlapped observed values of the two samples.
Hot-deck imputation is widely used due to its relative simplicity and its non-parametric nature, as it does not require assumptions about the underlying data distribution. However, the method has several limitations. It can underestimate data variability by redistributing existing values without introducing new information. Furthermore, imputation quality is highly dependent on the chosen similarity metric; a poor metric that fails to capture true associations between cases can lead to inaccurate imputations. This issue is exacerbated by high missingness rates or noisy data.
@@ 5.4 Likelihood Methods
Maximum likelihood is a strategy for estimating the unknown parameters of a model from data. When data is complete, it’s straightforward to calculate quantities like the mean, covariance, and linear regression coefficients, which are maximum likelihood estimates derived from maximizing the likelihood of the observed data. This same principle applies even when there are missing values in the data, as the estimation can still be based on the likelihood of the observed values. However, when missing data is present, the likelihood of the observed data becomes more complex compared to the typical data analysis scenarios, and finding an estimate that maximizes this likelihood doesn’t have a straightforward solution. To tackle the challenge, dempster1977maximum proposed the Expectation-Maximization (EM) algorithm for incomplete data. The main idea and algorithm are explained in the following context.
Following the notations defined by Section 3.1 and letting $\theta$ and $\phi$ be the parameters of the data model and the parameters of the missing mechanism, respectively, the full likelihood can be formulated as
@@
The two steps are repeated until some convergence conditions are met. Finally, the missing values are filled by using the estimated $\theta$. See the example of multivariate Gaussian in Section 11.6 of (murphy2012machine).
Despite its effectiveness, the EM algorithm has several major limitations. First, it relies on the MAR assumption and is not suitable for data that are MNAR. Second, its performance is contingent on a pre-specified data distribution (e.g., Gaussian); violations of this assumption can lead to inaccurate imputations. Finally, the algorithm is computationally intensive, and its results can be sensitive to the initial parameter values.
@@ 5.5.1 Low-Rank Matrix Completion
Low-rank matrix completion (LRMC) (srebro2005rank; CandesRecht2009; wen2012solving; hardt2014understanding) addresses the problem of recovering missing entries in a matrix under the assumption that the true matrix is low-rank. Formally, a matrix is low-rank if its rank (the dimension of the vector space spanned by its rows or columns) is significantly smaller than its dimensions. An equivalent characterization is that the matrix has a rapidly decaying singular value spectrum, with only a few non-zero or significant singular values. Consider a matrix $\mathbf{X}\in\mathbb{R}^{m\times n}$, where $m<n$, the singular value decomposition (SVD) is denoted as $\mathbf{X}=\mathbf{U}\bm{\Sigma}\mathbf{V}^{\top}$, where $\mathbf{U}\in\mathbb{R}^{m\times m}$ and $\mathbf{V}\in\mathbb{R}^{n\times n}$ are orthonormal matrices, i.e., $\mathbf{U}^{\top}\mathbf{U}=\mathbf{I}_{m}$ and $\mathbf{V}^{\top}\mathbf{V}=\mathbf{I}_{n}$, $\bm{\Sigma}=(\text{diag}(\sigma_{1},\sigma_{2},\ldots,\sigma_{m}),\bm{0})$, and $\sigma_{1}\geq\sigma_{2}\geq\cdots\geq\sigma_{m}$ are singular values. The rank of $\mathbf{X}$ is defined as $\text{rank}(\mathbf{X}):=\sum_{i=1}^{m}\mathbbm{1}(\sigma_{i}\neq 0)$. It has been found that many data matrices in real applications are low-rank ($\text{rank}(\mathbf{X})\ll m$) or can be well approximated by a low-rank matrix, i.e., many singular values are close to zero.
Over the past decades, numerous LRMC methods have been developed, which generally fall into two categories: low-rank regularization and low-rank factorization. Given an incomplete data matrix $\tilde{\mathbf{X}}\in\mathbb{R}^{m\times n}$, where the set of locations of observed entries is denoted as $\Omega$, the low-rank regularization methods aim to solve the following problem
@@
$\underset{\mathbf{A},\mathbf{B}}{\operatorname{minimize}}$ $\|\mathcal{P}_{\Omega}(\tilde{\mathbf{X}}-\mathbf{AB})\|_{F}^{2}+\lambda\cdot R(\mathbf{A},\mathbf{B})$ (11)
where $R$ denotes the regularization on the factors $\mathbf{A}$ and $\mathbf{B}$. Table 3 summarizes the choices for $R$, where most of them are closely related to the nuclear norm and Schatten-$p$ quasi-norms. These regularizers are more efficient to implement than those in Table 2 because they do not require SVD or only need to perform SVD on the smaller factors. Once $\mathbf{A}$ and $\mathbf{B}$ are found, the recovered matrix is $\hat{\mathbf{X}}=\mathbf{A}\mathbf{B}$.
@@
nuclear norm $\|\mathbf{X}\|_{\ast}=\sum_{i=1}^{m}\sigma_{i}(\mathbf{X})$ (CandesRecht2009)
truncated nuclear norm $\|\mathbf{X}\|_{\text{TN}}=\sum_{i=r+1}^{m}\sigma_{i}(\mathbf{X})$ (hu2012fast)
weighted nuclear norm $\|\mathbf{X}\|_{\text{WN}}=\sum_{i=1}^{m}w_{i}\sigma_{i}(\mathbf{X})$ (gu2014weighted)
Schatten-$p$ quasi-norm $\|\mathbf{X}\|_{\text{Sp}}=(\sum_{i=1}^{m}\sigma_{i}^{p}(\mathbf{X}))^{1/p}$ (MC_ShattenP_AAAI125165)
weighted Schatten-$p$ quasi-norm $\|\mathbf{X}\|_{\text{WSp}}=(\sum_{i=1}^{m}w_{i}\sigma_{i}^{p}(\mathbf{X}))^{1/p}$ (7539605)
generalized non-convex penalty $R(\mathbf{X})=\sum_{i=1}^{m}g_{\lambda}(\sigma_{i}(\mathbf{X}))$ (lu2014generalized)
@@
factored nuclear norm $R(\mathbf{A},\mathbf{B})=\frac{1}{2}(\|\mathbf{A}\|_{F}^{2}+\|\mathbf{B}\|_{F}^{2})$ $\min R(\mathbf{A},\mathbf{B})=\|\mathbf{X}\|_{\ast}$ (srebro2005maximum)
max norm $R(\mathbf{A},\mathbf{B})=(\max_{i}\left\|\mathbf{A}_{i:}\right\|_{2})(\max_{j}\left\|\mathbf{B}_{:j}\right\|_{2})$ srebro2005rank
factored Schatten-$1/2$ $R(\mathbf{A},\mathbf{B})=\frac{1}{2}(\|\mathbf{A}\|_{*}+\|\mathbf{B}\|_{*})$ $\min R^{2}(\mathbf{A},\mathbf{B})=\|\mathbf{X}\|_{\mathrm{S}_{1/2}}$ (shang2016tractable)
factored Schatten-$1/2$ $R(\mathbf{A},\mathbf{B})=\frac{1}{2}(\|\mathbf{A}\|_{2,1}+\left\|\mathbf{B}^{T}\right\|_{2,1})$ $\min R^{2}(\mathbf{A},\mathbf{B})=\|\mathbf{X}\|_{\mathrm{S}_{1/2}}$ (fan2019factor)
factored Schatten-$2q/(2+q)$ $R(\mathbf{A},\mathbf{B})=\frac{1}{q}\|\mathbf{A}\|_{2,q}^{q}+\frac{\alpha}{2}\left\|\mathbf{B}^{T}\right\|_{F}^{2}$ $\min R(\mathbf{A},\mathbf{B})=c\|\mathbf{X}\|_{S_{2q/(2+q)}}^{2q/(2+q)}$ (fan2019factor)
factored Schatten-$p$ $R(\mathbf{A},\mathbf{B})=\frac{1}{2^{p}}\sum_{i=1}^{d}\left(\left\|\mathbf{A}_{:i}\right\|_{2}^{2}+\left\|\mathbf{B}_{i:}\right\|_{2}^{2}\right)^{p}$ $\min R(\mathbf{A},\mathbf{B})=\|\mathbf{X}\|_{S_{p}}^{p}$ (giampouras2020novel)
@@
Beyond general low-rank matrices, several LRMC methods are designed for specific types. Notable examples include distance matrix completion (alfakih1999solving), covariance matrix completion (hippert2022robust), kernel matrix completion (tsuda2003algorithm), and one-bit matrix completion (davenport20141). In the latter case, the observed data are binary (1-bit) measurements rather than precise numerical values. This approach is valuable in scenarios with highly quantized data or binary feedback, such as in recommender systems and social networks.
@@ 5.5.2 High-Rank Matrix Completion
where $f^{(1)},\ldots,f^{(s)}$ are different nonlinear functions. This forms a matrix $\mathbf{X}=[\mathbf{X}^{(1)},\mathbf{X}^{(2)},\ldots,\mathbf{X}^{(s)}]$ of size $m\times n$, where $\mathbf{X}^{(j)}=[\mathbf{x}_{1}^{(j)},\ldots,\mathbf{x}^{(j)}_{n_{j}}]\in\mathbb{R}^{m\times n_{j}}$ and $n=\sum_{j=1}^{s}n_{j}$. It can be seen that (14) degrades to (13) if $s=1$ and degrades to (12) if all $f^{(j)}$ are linear. Figure 8 provides synthetic examples for (12),(13), and (14), respectively. The completion of the matrix produced by (14) poses the greatest challenge.
Figure 8: Examples (fan2020polynomial) of data forming high-rank matrices in 3D space (left: union of subspaces; middle: one nonlinear manifold; right: union of nonlinear manifolds).
In recent years, a few high-rank matrix completion (HRMC) methods have been proposed (ErikssonBalzanoNowak2011; NLMC2016; NIPS2016_6357; fan2018matrix; fan2020polynomial; NEURIPS2020_42ae1544). They can be organized into three categories: subspace-based methods, kernel-based methods, and deep learning-based methods.
@@
These methods were proposed for the data given by (12) and usually learn the subspaces from the incomplete data (ErikssonBalzanoNowak2011; 8469061) or exploit the self-expressive model (yang2015sparse; NIPS2016_6357; fan2017sparse; FAN2017290), where each data point can be well represented by other data points drawn from the same subspace. The self-expressive model (elhamifar2013sparse) can be formulated as
@@
where $\mathbf{S}\in\mathbb{R}^{n\times n}$ is the coefficient matrix and is usually assumed to be sparse. Given an incomplete data matrix $\tilde{\mathbf{X}}$, FAN2017290 proposed the following matrix completion method:
@@
[…]
@@
These methods (NLMC2016; pmlr-v70-ongie17a; FANNLMC; fan2019online; fan2020polynomial; ongie2021tensor) take advantage of kernel techniques that can convert a nonlinear problem in the original data space to a linear problem in the high-dimensional feature space. pmlr-v70-ongie17a and FANNLMC proposed to solve
@@
[…]
fan2019online showed that $\text{rank}(\phi(\mathbf{X}))=\min\{\binom{m+q}{q},n,s\binom{d+pq}{pq}\}$ if $f^{(j)}$ are $p$-order polynomials and $\phi$ is the feature map of a $q$-order polynomial kernel. When $n$ is sufficiently large and $p$ and $q$ are relatively small, the rank of $\phi(\mathbf{X})$ is low compared to its dimensions. fan2019online introduced a kernel factorization method for HRMC and an online algorithm, which are more efficient than the methods proposed in (pmlr-v70-ongie17a; FANNLMC).
@@
These methods (FAN2017540; fan2018matrix; seyedi2023elastic) are based on neural networks and will be detailed in Section 5.6.